# The Lazy Student’s Dream: ChatGPT Passing an Engineering Course on Its Own

### v2

@@ 1 Introduction
Statement of contributions: Our research deliberately maintains a focused scope, examining LLM performance in a single undergraduate control systems course rather than attempting to draw broad conclusions about AI in education. The significance of this work lies in three key areas: First, we establish a rigorous, reproducible methodology for evaluating LLM performance across diverse assessment types within technical courses. Second, we provide quantitative insights into the current capabilities and limitations of LLMs in completing complex engineering coursework.
@@ 2.3 AE 353: Course Structure and Requirements
Three substantial programming projects represent the course’s practical implementation component. These projects demand extensive Python coding, data analysis, and technical writing skills. Each project builds upon theoretical concepts covered in lectures, requiring students to implement control algorithms, analyze system behavior, and document their findings in detailed technical reports.
@@ Acknowledgments
This work was supported by the Grants for Advancement of Teaching in Engineering program at the Grainger College of Engineering, University of Illinois Urbana-Champaign. We gratefully acknowledge Prof. Timothy Bretl for providing the course materials, assignments, projects, and PrairieLearn infrastructure foundational to this study. We also thank Grayson Schaer for developing the project environments and Pranay Thangeda for contributing to the PrairieLearn question bank and lecture materials. Additionally, we extend appreciation to all others whose efforts supported this research.

### v4

@@ 1 Introduction
Statement of contributions: Our research deliberately maintains a focused scope, examining LLM performance in a single undergraduate control systems course rather than attempting to draw broad conclusions about AI in education. The significance of this work lies in two key areas: First, we establish a rigorous, reproducible methodology for evaluating LLM performance across diverse assessment types within technical courses. Second, we provide quantitative insights into the current capabilities and limitations of LLMs in completing complex engineering coursework.
@@ 2.3 AE 353: Course Structure and Requirements
Three substantial programming projects represent the course’s practical implementation component. These proje-cts demand extensive Python coding, data analysis, and technical writing skills. Each project builds upon theoretical concepts covered in lectures, requiring students to implement control algorithms, analyze system behavior, and document their findings in detailed technical reports.
@@ Acknowledgments
This study was supported by the University of Illinois Urbana-Champaign’s Grainger College of Engineering through its Grants for the Advancement of Teaching in Engineering program. The course’s question bank represents several years of iterative development by the authors. We gratefully acknowledge Grayson Schaer for building the project environments, and David Hanley and Pranay Thangeda for their significant contributions to the PrairieLearn question bank and lecture materials. We also thank the many others whose efforts helped make this research possible.